You need to call a plumber to fix the sink in your kitchen. The first company you call gives a price of $60 for the house-call, plus $25 per hour. The second company charges $45 per hour, with no extra charge for the house-call. After how many hours would the two costs be equal?
Answer by checkley79(3341)

60+25X=45X
25X-45X=-60
-20X=-60
X=-60/-20
X=3 HOURS ANS.
PROOF:
60+25*3=45*3
60+75=135
135=135